OpenStudy (anonymous):

I will give an medal or possibly 2 for answering these questions [9.01] Determine whether the graph of y = x2 − 4x + 1 has a maximum or minimum point, then find the maximum or minimum value. Minimum; (2, -3) Maximum; (2, -3) Minimum; (-3, 2) Maximum; (-3, 2) 7. [9.02] What are the x-intercept(s) of the graph of y − 8 = x2 − 6x? (−2, 0) and (4, 0) (2, 0) and (−4, 0) (2, 0) and (4, 0) (−2, 0) and (−4, 0) 8. [9.02] What are the solutions of 5x2 = −15x? x = 5, x = 3 x = 5, x = −3 x = 0, x = 3 x = 0, x = −3 9. [9.02] Where does the graph of y = 2x2 − 5x + 3 cross the x-axis?

OpenStudy (anonymous):

[9.02] Where does the graph of y = 2x2 − 5x + 3 cross the x-axis? (−1 over 2, 0) and (−3, 0) (1 over 2, 0) and (3, 0) (3 over 2, 0) and (1, 0) (−3 over 2, 0) and (−1, 0) 10. [9.02] Solve x2 − 3x = 28. x = −2, x = 14 x = 2, x = −14 x = 4, x = −7 x = −4, x = 7

OpenStudy (anonymous):

satellite how did i know you would be back lol

OpenStudy (anonymous):

just unlucky i guess

OpenStudy (anonymous):

\[y = x^2 − 4x + 1\] leading coefficient is positive, so it opens up and has a minimum

OpenStudy (anonymous):

ok then but if i did my math right should it be minumum (2,-3)?

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is 2

OpenStudy (anonymous):

second is what you get, which is suppose is -3 but the question as stated is in fact incorrect

OpenStudy (anonymous):

so wait now im confused is minimum (2,-3) not right then?

OpenStudy (anonymous):

that is what you are supposed to say, yes

OpenStudy (anonymous):

oh ok then what about the rest

OpenStudy (whpalmer4):

what is incorrect about the question?

OpenStudy (anonymous):

but the minimum value is a number, it is \(-3\) the ordered pair is the vertex

OpenStudy (whpalmer4):

minimum point is what it says

OpenStudy (anonymous):

if i ask you what the least amount of money you would pay me, you would give a number as an answer, not an ordered pair this question was written by someone who is guilty of causing you confusion later on no matter, just give them the answer they want

OpenStudy (anonymous):

there is no such thing as a "minimum point"

OpenStudy (anonymous):

Well im about done with this algebra thats why im trying to get help with these cause my final will be coming soon

OpenStudy (anonymous):

good

OpenStudy (anonymous):

well not soon but in december

OpenStudy (anonymous):

\[ y − 8 = x^2 − 6x\] is the same as \[y=x^2-6x+8\] the \(x\) intercepts you find by setting \[x^2-6x+8=0\] and solving for \(x\)

OpenStudy (anonymous):

factor as \((x-2)(x-4)=0\) and the it should be easy to solve

OpenStudy (anonymous):

yea ok i got 7 now but what about the rest of the questions?

OpenStudy (anonymous):

\\[5x^2 = −15x\] is the same as \[x^2=-3x\] rewrite as \(x^2+3x=0\) and the factor as \(x(x+3)=0\) and solve that one

OpenStudy (anonymous):

then if its 0, then i would think it should be -3? am i right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

woohoo im learning

OpenStudy (anonymous):

two solutions \(x=0\) or \(x=-3\)

OpenStudy (anonymous):

ok, but 9 really confuses me

OpenStudy (anonymous):

\[y = 2x^2 − 5x + 3\] again set \[2x^2-5x+3=0\] and solve by factoring

OpenStudy (anonymous):

try \[(x-1)(2x-3)=0\]

OpenStudy (anonymous):

so, if kinda confused but im thinking it might be something like (3,0) for one half of the answer

OpenStudy (anonymous):

no not \((3,0)\) solve the one i wrote above

OpenStudy (anonymous):

oh ok, sure like i said confusing i dont even know basically where to start but if i did it right would it be something like 3x-4 = 0? or am i not even close

OpenStudy (anonymous):

then like 3x = 4 then divide into a fraction?

OpenStudy (anonymous):

do you know how to solve \[(x-1)(2x-3)=0\] for \(x\) ?

OpenStudy (anonymous):

not exactly

OpenStudy (anonymous):

set \(x-1=0\) first

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

just add the one to the other side?

OpenStudy (anonymous):

if \(x-1=0\) then \(x=1\) right

OpenStudy (anonymous):

ok then what did i do wrong then in the equation you gave me

OpenStudy (anonymous):

and also set \(2x-3=0\) and solve

OpenStudy (anonymous):

2x = 3 then it would work out into to be 3 over 2 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so the first half is (3/2, 0)

OpenStudy (anonymous):

ok i got it now i got 9 for the final answer to be (3/2,0) and (1,0)

OpenStudy (anonymous):

I only need 10 now